OpenStudy (anonymous):

two rubber balls collide in a collision, one ball has a mass of 2 kg and has a velocity of 6m/s due east. the other ball has a mass of 6 kg and is traveling behind the other ball at a velocity of 8 m/sec due east. after the collision, the 6 kg ball slows down to 7 m/sec . find the velocity of the 2 kg ball after the collision

OpenStudy (anonymous):

use momentum conservation mv+mv=mv+mv 2(6)+6(8)=6(7)+2(x) Solve for x

OpenStudy (anonymous):

thanks @cherio12